OpenStudy (anonymous):

derivative of 4cos[5pi(x)] sorry for my terrible syntax..

OpenStudy (anonymous):

\[-4\sin(5\pi x)\times 5\pi=-20\pi\sin(5\pi x)\]

OpenStudy (anonymous):

but how i got lost in my u-sub

OpenStudy (anonymous):

you integrating or differentiating?

OpenStudy (anonymous):

diff right? then by the chain rule \[\frac{d}{dx}[cos(ax)]=-\sin(ax)\times a=-a\sin(ax)\]

OpenStudy (anonymous):

oh okay.. wolframalpha was confusing me thank you

OpenStudy (anonymous):

you can write \(u=5\pi x, u'=5\pi\) if you like but it is a waste of time, you should do it in your head derivative of the outside function evaluated at the inside function time the derivative of the inside function

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

yw